# Did I Faithfully Say What I Thought? Bridging the Gap Between Neural Activity and Self-Explanations in Large Language Models

### v1

@@ 5.1 Datasets and Models
We apply NeuroFaith on the Wikidata-2-hop dataset Biran et al. (2024). Following the notations introduced in Section 4, each input text is formatted as $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$. We evaluate Gemma-2-2B and Gemma-2-9B Riviere et al. (2024) in their instruction-tuned versions, using the following backbones from the transformers library Wolf et al. (2020): gemma-2-2B-it and gemma-2-9B-it.
@@ 5.2 Framework Instantiation
    Hidden States Interpretation. We use Selfie Chen et al. (2024) as interpreter $\mathcal{I}$ to decode the $f$ hidden states of interest and generate a set of local neural interpretations $i(x)$. We focus on the residual stream ($\mathcal{L}=\texttt{RS}$) and define the circuit $\mathcal{C}$ based on recent work Biran et al. (2024); Yang et al. (2024), which identifies the latent first hop operation in the early layers (when processing the last source token corresponding $o_{1}$). We therefore set the index token $k$ as the last one related to $o_{1}$ and $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B-it and $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B-it.
@@ 6 Discussion
In this paper, we focus on generating free text interpretations for faithfulness assessment. Future work could explore additional NeuroFaith implementations, such as using Sparse Autoencoders (SAE) to generate concept-based interpretations. Additionally, our detailed characterization focused exclusively on cases with correct final predictions ($\widehat{o_{3}}=o_{3}$). Developing similar characterizations when models make incorrect predictions would provide valuable insights. Our current study calculated faithfulness of free text self-NLE only in predict-then-explain contexts. A promising extension would be to compare the faithfulness of self-NLE with the performance gains they can induce Bhan et al. (2024). We believe that NeuroFaith could also lead to valuable results by measuring CoT faithfulness and being compared to other studies in this direction Lanham et al. (2023); Turpin et al. (2023). Examining the faithfulness of alternative explanation methods such as counterfactuals Madsen et al. (2024) or concept-based approaches Bhan et al. (2025) through comparison with the model’s internal neural activity is another promising research direction.

### v2

@@ 5.1 Datasets and Models
We apply NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024). Following the notations introduced in Section 4, each input text is formatted as $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$. We evaluate Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024) in their instruction-tuned versions, using the following backbones from the transformers library (Wolf et al., 2020): gemma-2-2B-it and gemma-2-9B-it.
@@ 5.2 Framework Instantiation
    Hidden States Interpretation. We use Selfie (Chen et al., 2024) as interpreter $\mathcal{I}$ to decode the $f$ hidden states of interest and generate a set of local neural interpretations $i(x)$. We focus on the residual stream ($\mathcal{L}=\texttt{RS}$) and define the circuit $\mathcal{C}$ based on recent work (Biran et al., 2024; Yang et al., 2024), which identifies the latent first hop operation in the early layers (when processing the last source token corresponding $o_{1}$). We therefore set the index token $k$ as the last one related to $o_{1}$ and $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B-it and $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B-it.
@@ 6 Discussion
In this paper, we focus on generating free text interpretations for faithfulness assessment. Future work could explore additional NeuroFaith implementations, such as using Sparse Autoencoders (SAE) to generate concept-based interpretations. Additionally, our detailed characterization focused exclusively on cases with correct final predictions ($\widehat{o_{3}}=o_{3}$). Developing similar characterizations when models make incorrect predictions would provide valuable insights. Our current study calculated faithfulness of free text self-NLE only in predict-then-explain contexts. A promising extension would be to compare the faithfulness of self-NLE with the performance gains they can induce (Bhan et al., 2024). We believe that NeuroFaith could also lead to valuable results by measuring CoT faithfulness and being compared to other studies in this direction (Lanham et al., 2023; Turpin et al., 2023). Examining the faithfulness of alternative explanation methods such as counterfactuals (Madsen et al., 2024) or concept-based approaches (Bhan et al., 2025) through comparison with the model’s internal neural activity is another promising research direction.